OpenStudy (anonymous):

f(x)=x^4+2x^3-4x^2-x+1 let a,b,c,d be the roots of this ab, ac, ad, bc, bd, cd are roots of polynomial g(x) what is g(1) if g(0)=1

OpenStudy (anonymous):

wow i am stumped, must be something to do with vieta's formulas, elementary symmetric functions g(1) is the sum of the coefficients, so that is what you are being asked for, but i am not sure how to get them. do you have any clues?

OpenStudy (anonymous):

when you add these together you should get \[\frac{a_2}{a_4}=\frac{-4}{1}=-4\]

OpenStudy (anonymous):

and also \[a+b+c+d=-\frac{a_3}{a_4}=-2\] if that helps

OpenStudy (anonymous):

i will have to think on this, converting roots of one poly to roots of another, but if you get an answer please post, i would love to see it

OpenStudy (anonymous):

I've been pondering with this the entire day. The problem asks us to find g(1).

OpenStudy (anonymous):

this is what we know about a, b , c, d from vieta, \[a+b+c+d=-2\] \[ab+ac+ad+bc+bd+cd=-4\] \[abc+abd+acd+bcd=1\] \[abcd=1\]

OpenStudy (anonymous):

we also know that we have a sixth degree polynomial that looks like \[(x-ab)(x-ac)(x-ad)(x-bc)(x-bd)(x-cd)\]

OpenStudy (anonymous):

and we are being asked for the sum of the coeffidents, because that is what \[g(1)\] gives

OpenStudy (anonymous):

Let a,b,c, and d be distinct roots of \[f(x)=x^4+2x^3-4x^2-x+1\]The six values ab,ac,ad,bc,bd and cd are roots of the polynomial \[g(x)\] If \[g(0)=1\] find \[g(1)\]

OpenStudy (anonymous):

If that helps

OpenStudy (anonymous):

ok maybe we are getting somewhere

OpenStudy (anonymous):

no i got the question

OpenStudy (anonymous):

you are being asked for the sum of the coefficents of g

OpenStudy (anonymous):

i think this might be hard, but i am not sure. where did the problem come from?

OpenStudy (anonymous):

I was working on it with a teacher for fun the other day, neither of us found an answer, even as he was looking up theorems to help solve it

OpenStudy (anonymous):

i think you need vieta's formula, the one that gives the coefficients of a polynomial in terms of the sum and the product of the roots. you can look here http://en.wikipedia.org/wiki/Vieta%27s_formulas or in this nice pdf attachment on page 68

OpenStudy (anonymous):

keep in mind that g(1) is the sum of the coefficents, so that is what you are being asked for, and g(0)=1 tells you that the product of the roots \[a^3b^3c^3d^3=1\]

OpenStudy (anonymous):

Alright, I'll try to work off that for the rest of the night. Thank you very much!

OpenStudy (anonymous):

i am afraid i was not much help but it might be instructive to do a simple example, one with the roots that you know

OpenStudy (anonymous):

say \[x^4-14 x^3+71 x^2-154 x+120=(x-2)(x-3)(x-4)(x-5)\] then take the product of the roots taken two at a time and see how the coefficients relate

OpenStudy (anonymous):

\[(x-6)(x-8)(x-10)(x-12)(x-15)(x-20)=x^6-71 x^5+2036 x^4-30196 x^3+244320 x^2-1022400 x+1728000\]

OpenStudy (anonymous):

sum of coefficients is 921690

OpenStudy (anonymous):

ok so much for that bright idea, i don't see the connection, but who knows. i will post also so maybe tomorrow someone will give a good explanation

OpenStudy (anonymous):

Thank you very much